OpenStudy (anonymous):

A particle moves along a straight line such that its displacement at any time t is given by S= t 3 - 3t 2 + 2 meter. The displacement when the acceleration becomes zero is

OpenStudy (badhi):

acceleration can be taken by differentiating S twice wrt t ie. $$S=t^3-3t^2+2\implies \frac{dS}{dt}=3t^2-6t\implies \frac{d^2S}{dt^2}=6t-6$$ when acceleration becomes zero, $$6t-6=0\implies t=1$$ when t=1, $$S=1-3+2=0$$

OpenStudy (anonymous):

Thnx Badhi